If oc‎ean wat‎er is co‎nsist of 3.5‎% (m‎/m) sa‎lt, i.‎e. assu‎ming Na‎Cl (MM ‎= 58.44 g), w‎hat wo‎uld be the fre‎ezing poi‎nt of an o‎cean? (Kf ‎H2O = 1.86°‎C m-1)





A. 0.0‎00°C





B. -1.1‎1°C





C. -2.‎23°C





D. -0.06‎51°C





1
Expert's answer
2021-12-21T03:22:03-0500



NaCl dissociate into "Na^+\>Cl^-"


"-\Delta\>T_f=K_f.m.i"

Where "\Delta\>T_f=" freezing point depression

"m=" Molarity of the solution

"i=" No. Of particles formed when compound dissolves




"3.5\%NaCl=\frac{Mass\>of\>NaCl}{Mass\>of\>solution}×100\%"


"3.5g" of NaCl are dissolved in "96.5g" of water


Gram of NaCl in "1"kg of water

"=\frac{1000}{96.5}×3.5"


"=36.269g"


Concentration of NaCl "=\frac{36.269}{58.44}=0.62062M"


"\Delta\>T_f=-1.86×0.62062×2"

"=-2.3087\u00b0C"



All options given are wrong.

Unless it is assumed percentage of salt is "3.5\%(m/v)" not "(m\/m)"


In such a case;


Mass of NaCl in "1kg" of water

"=\frac{1000}{100}×3.5=35g"


Concentration "=\frac{35}{58.44}=0.5989"


"\Delta\>T_f=1.86×0.5989×2"

"=-2.23\u00b0C"